OpenStudy (anonymous):

1) y ={ 1/x+1 } +4 2) 2x+8/3x-9 Domain Range V Asym H Asym Y-Intercept Graph:

OpenStudy (anonymous):

I need help here @goformit100

OpenStudy (goformit100):

Madam, sorry but can you wait for me, I am solving it.

OpenStudy (anonymous):

Sure no problem

OpenStudy (anonymous):

Is that for the first 1? And I need to fill out the domain & range below .

OpenStudy (goformit100):

It was for the second one.

OpenStudy (anonymous):

Okay thank you. Domain Range V Asym H Asym Y-Intercept

OpenStudy (goformit100):

Asym will be denoting the line satisfying and touching the Maximum point of the graph

OpenStudy (anonymous):

There are two possibilities because you didn't really specify the problem. Either Y = 1/(X+1) OR Y = 1/X + 1 Either way, firs tyou substitute 4 in for Y in the first equation. For the first possibility 4 = 1/(X+1) 4(X+1) = 1 (Multiply both sides by X+1) 4X + 4 = 1 (Distribute) 4X = -3 (Subtract 4) X = -3/4 (Divide by 4) Second possibility 4 = 1/X +1 3 = 1/X (Subtract 1) 3X = 1 (Multiply by X) X = 1/3 (Divide by 3)

OpenStudy (anonymous):

For number one right?

OpenStudy (anonymous):

Oh it's not solving for x . It's graphing the equation. @steffanie1995

OpenStudy (anonymous):

ohh my bad.. Hold up

OpenStudy (goformit100):

@steffanie1995 yes madam plz go through the question again.

OpenStudy (anonymous):

Sorry I don't know. I'm bad at math.

OpenStudy (anonymous):

Haha no problem as well am I.

OpenStudy (goformit100):

Domain for both the equations will be Real Numbers

OpenStudy (goformit100):

Did you understood till here. ?

OpenStudy (anonymous):

Hmmm a little bit.

OpenStudy (goformit100):

And what left to be unanswered Plz Mention it madam.

OpenStudy (mathstudent55):

1. Domain: What value makes denominator zero? Solve x + 1 = 0 That value must be excluded from domain.

OpenStudy (anonymous):

For the second graph I have figured out: D: \[x \neq \frac{ 2 }{ 3 }\] R: \[y \neq 3\] V Asym: 3 H Asym: 2/3 Y Intercept : ?

OpenStudy (goformit100):

Your above solutions are correct

OpenStudy (goformit100):

Now its the matter of Y...

OpenStudy (anonymous):

@mathstudent55 Is the above correct?

OpenStudy (goformit100):

Yes above one is correct

OpenStudy (goformit100):

Hey its correct, no worries @Ozzy733

sam (.sam.):

For y-intercept just think of x=0, so \[y=\frac{2x+8}{3x-9} \\ \\ y=\frac{2(0)+8}{3(0)-9} \\ \\ y=-\frac{8}{9}\] \[\large (0,-\frac{8}{9})\]

OpenStudy (anonymous):

Thanks! While I solve the y-intercept for the first graph, could you tell me what the h asym and domain would be for the first graph? As well as the graph sketch.

OpenStudy (anonymous):

Would the y-intercept for the first graph just be 5?

OpenStudy (anonymous):

@.Sam. ^

sam (.sam.):

For the first one is it\[y=\frac{1}{x+1}+4\]

OpenStudy (goformit100):

@Ozzy733 @.Sam.

OpenStudy (anonymous):

Yes it is

sam (.sam.):

It's a reciprocal with the graph shifted upwards by 4

sam (.sam.):

yes the y-intercept is 5

OpenStudy (anonymous):

Okay great! I'm stuck on the domain & hasym. I know it's a/c but I don't understand how to get it from this setup.

sam (.sam.):

For Horizontal asymptote, as "x" approaches -1, the equation becomes infinity For Vertical asymptote, as "x" approaches infinity, the \(\huge\frac{1}{x+1}\) becomes too small and we can just call it zero, so it leaves with 4.

OpenStudy (goformit100):

My Doubt :- Synthetic polymers can be divided into two major classes, depending on their method of preparation. Chain-growth polymers, also known as addition polymers, are made by chain reactions—the addition of monomers to the end of a growing chain. The end of the chain is reactive because it is a radical, a cation, or an anion. Polystyrene—used for disposable food containers, insulation, and toothbrush handles, among other things—is an example of a chain-growth polymer. Polystyrene is pumped full of air to produce the material known as Styrofoam®.

OpenStudy (anonymous):

So the asym is 0?

OpenStudy (anonymous):

Thanks so much!

sam (.sam.):

H:-1 V:4

sam (.sam.):

welcome :)

OpenStudy (anonymous):

So then the domain would be \[x \neq 4\] correct?

OpenStudy (goformit100):

Yes the domain would be x≠4 is correct.